Refer to the data below which contains the percentage change​ (X) in a stock market index over the first five trading days of the year and percentage change​ (Y) in the index over the whole year. Complete parts a and b.

​(X)

1.4

0.3

−0.2

2.9

2.2

−1.9

−1.5

5.4

−1.6

1.4

1.7

−4.8

−0.1

​(Y)

14.9

−9.4

19.4

20.3

−3.8

27.6

22.4

2.4

11.9

28

−4.4

20.5

−0.3

a. Estimate the linear regression of Y on X.

Yi=_______ +(______)Xi

​(Round to four decimal places as​ needed.)

b. Provide interpretations of the intercept and slope of the sample regression line.

For a one percent gain or loss ? over the first five trading days of the​ year, estimates would predict a change of ______%  over the whole year. If there is a zero percent change over the first five trading days of the​ year, estimates would predict a change of ___________​% over the whole year.
Answer #1

a. Х- М., V.MN (Х- Му)? (Х- М.ХҮ-М) е н о л со om in m л ооо оо л со уоu са о о л вно носа о о в е уол муо: UwNNO ш о л н м -15.

Sum of X = 5.2
Sum of Y = 149.5
Mean X = 0.4
Mean Y = 11.5
Sum of squares (SSX) = 78.74
Sum of products (SP) = -153.9

Regression Equation = ŷ = bX + a

b = SP/SSX = -153.9/78.74 = -1.9545

a = MY - bMX = 11.5 - (-1.95*0.4) = 12.2818

ŷ = -1.9545X + 12.2818

b. For a one percent loss over the first five trading days of the​ year, estimates would predict a change of -1.9545%  over the whole year.

If there is a zero percent change over the first five trading days of the​ year, estimates would predict a change of 12.2818​% over the whole year.